Integrate (1/(x+1)) dx within the limits 0 to 4 in four parts using Simpson rule


1
Expert's answer
2022-04-16T04:13:06-0400

Solution

For function f(x) with division interval [a,b] in four parts

h = (b-a)/4, x0 = a, x1 = a+h, x2 = a+2h, x3 = a+3h, x4 = b

For given integral

h = 1, x0 = 0, x1 = 1, x2 = 2, x3 = 3, x4 = 4

If fi = f(xi) = 1/(xi+1) then integral value using Simpson rule is

I = (f0+4f1+2f2+4f3+f4)/3

f0 = 1, f1 = 0.5, f2 = 0.33333, f3 = 0.25, f4 = 0.2

Therefore I = 1.622

Exact value of given integral is ln(5) = 1.609

Solution I = 1.622